# Fast and Accurate Algorithms to Calculate Expected Modularity in Probabilistic Networks

### v2

@@ 4.3 Factors influencing execution time
In Figure 7, we plot the execution time of $\mathrm{APWP}^{\mathrm{EMOD}}$ depending on the number of communities. For this experiment we generated a random probabilistic graph with 100 nodes and 500 edges. We vary the number of communities from 3 to 20, randomly assigning the nodes to them. From this figure, we observe that as the number of communities increases, the execution time calculated by $\mathrm{APWP}^{\mathrm{EMOD}}$ also lengthens.
@@
In Figure 8, we show the performance of $\mathrm{APWP}^{\mathrm{EMOD}}$ with the same number of communities but different community size distributions. From this figure, we can find that the larger variance among communities, the longer the running time.
@@
Figure 7: Time calculated by $\mathrm{APWP}^{\mathrm{EMOD}}$ with different numbers of communities.
@@
Figure 8: Time calculated by $\mathrm{APWP}^{\mathrm{EMOD}}$ with different variances of communities size.
@@
In our final experiment, we examine the execution time of $\mathrm{APWP}^{\mathrm{EMOD}}$ on different random networks. In particular, we use 5 random graph models: a Forest Fire Network (FFN) [28], Barabási-Abert (BA) [29], Small world (SW) [30], Erdős-Rényi (ER) [31], and a clear community-structure graph (CCS graph). We use each of these models to generate three random graphs, all with around 200 nodes and 600 edges, but different numbers of communities. To provide a community assignment for the computation of modularity, we use a modularity optimization on the original deterministic graphs. We also control the number of communities, where $|C|=4$, $|C|=5$ and $|C|=6$ respectively.
@@
Figure 10 shows the running time of $\mathrm{APWP}^{\mathrm{EMOD}}$ on graphs with different network structures and numbers of communities. Generally, with an increasing number of communities the running time also increases, which fits the results in Figure 7. The visualization in Figure 9 suggests that a modularity optimization method may find a larger central community in FFN, which explains why the running time of FFN is always longer than for other networks. In fact, as shown in Figure 8, the larger the community size variance, the longer the running time.
@@
4.4 Real data
In this section, we apply $\mathrm{APWP}^{\mathrm{EMOD}}$ and the sampling method on real networks. Notice that we do not use thresholding and weighting, because the experiments on synthetic data have clearly shown that these approaches should not be used to compute expected modularity. Figure 12 shows different network structures with different numbers of clusters. Here nodes of the same color are in the same community computed using a modularity maximization method, and the number of colors represents the number of communities. Figure 11 shows the execution time (vertical blue line) and the computed expected modularity (blue horizontal line) for our algorithm. Each red circle is the result of a different execution of the sampling method. For different executions, we have let the sampling method run for different amounts of time, so that we could inspect the balance between execution time and the likelihood of returning an accurate result for the sampling method. Figure 11 shows the same trends seen in the experiments with synthetic data. In these datasets, where the probability distributions are fixed and where clear community structures exist (we remind the reader that real data was obtained from high-modularity networks), the sampling method can produce accurate results in a short time.
@@ 4.4 Real data
(e) Collaboration network $|E|=1224,|C|=2$
(f) PPI network $|E|=1185,|C|=5$
(g) Enron network $|E|=833,|C|=3$
(h) Facebook network $|E|=1217,|C|=4$
@@
5 Discussion
In this paper, we present two methods for uncertain graph expected modularity calculation based on the partitioning of possible worlds.
@@ 5 Discussion
When compared with a brute-force approach, both our methods are several orders of magnitude faster, with the approximate method reducing time complexity from exponential to polynomial without significantly affecting accuracy. For a network with 25 edges, which is close to the limit of what we could handle using exact methods before this paper, the execution time goes down from more than 15 hours (brute-force) to less than 5 seconds (our exact method) to a few microseconds (our DFT-based method). While this enables the usage of expected modularity in other algorithms, e.g. to identify communities, the execution time is however still high for large networks if the algorithm has to be repeated multiple times sequentially, taking several minutes for networks with thousands of edges.
@@
Our method is in general preferable to other common approaches used to handle probabilistic networks. Weighting should not be used as it returns wrong results. The main reason is that node strengths (which equal node expected degrees in probabilistic networks) are multiplied in weighted modularity computation, while the expected product of degrees is not the same as the product of expected degrees in case of dependencies. Unfortunately, degrees in probabilistic networks are not independent, even when edge probabilities are, because edges are incident to two nodes and thus they either contribute to the degree of both or of none of them. A second reason is that weights are interpreted relative to each other in weighted modularity. So, for example, having all weights set to 0.01 or having them set to 1 makes no difference, while interpreting these numbers as probabilities means that in one case the network is probably empty, while in the second case all edges are almost certainly there.
@@
In very specific cases, thresholding can be a fast alternative to our methods. However this is the case only when the network is almost deterministic. If we assume that all probabilities are at distance $\epsilon$ from 0 or 1, using thresholding we change edge probabilities from $1-\epsilon$ to 1 and from $\epsilon$ to 0, for all thresholds between $\epsilon$ and $1-\epsilon$. With small values of $\epsilon$, the change is small and the range of good thresholds is wide, making the method fast, accurate, and robust. However, as we are interested in handling probabilistic networks, it is not very useful to have a method that only works when the network is practically not probabilistic.
@@
A similar consideration holds for sampling: sampling from a probabilistic graph with probabilities very close to 0 or 1 immediately gives an accurate result, the same computed on the thresholded deterministic network. We have also noticed how the likelihood of producing an accurate modularity using sampling increases when the community structure is very clear, close to a set of cliques. This is due to the fact that modularity depends on the number of edges inside communities, and not on the position of these edges. As a result, when many (and most) edges are inside communities, many sampled graphs will have a similar number of edges inside those communities, and thus similar values of modularity. Sampling from a narrow distribution has then a faster convergence. However, in most interesting cases our algorithm is preferable to sampling: it returns an accurate value whose accuracy does not depend on the execution time, while sampling can give an inaccurate result, in addition to not giving any indication of whether the result is accurate or not. The more the network is uncertain, the less well sampling works, while our algorithms are not affected by the probability distribution (neither with respect to accuracy nor execution time).
@@
This paper opens the problem of how to use our algorithm as a sub-procedure of a community detection method. This problem adds two interesting aspects to our work: first, when used as a heuristic inside an optimization algorithm, a less accurate calculation of modularity may still lead to the same clustering, so there is a question about when it is needed to be accurate and when a faster approximation is sufficient. Second, we may have to execute the modularity calculation many times, which poses an additional computational challenge in particular for large networks. These two aspects may allow the usage of different computational approaches, for example to optimize a sequence of executions utilizing methods similar to the ones used in [32, 33].
@@ Acknowledgments
This work has been partly funded by eSSENCE, an e-Science collaboration funded as a strategic research area of Sweden, by the Centre of Natural Hazards and Disaster Science, by the Wallenberg AI, Autonomous Systems and Software Program WASP, and by the AI4Research initiative at Uppsala University.

### v3

@@ 17                         $Sum=Sum+Q_{c}^{xyz}P_{c}[x]P_{c,\bar{c}}[y]P_{\bar{c}}[z]$;
In Figure 9, we plot the execution time of $\mathrm{FPWP}$ depending on the number of communities. For this experiment we generated a random probabilistic network with 100 nodes and 500 edges. We vary the number of communities from 3 to 20, randomly assigning the nodes to them. From this figure, we observe that as the number of communities increases, the execution time of $\mathrm{FPWP}$ also increases.
@@
In Figure 10, we show the performance of $\mathrm{FPWP}$ with the same number of communities but different community size distributions. From this figure, we can find that the larger variance among communities, the longer the running time.
@@
Figure 9: Time calculated by $\mathrm{FPWP}$ with different numbers of communities.
@@
Figure 10: Time calculated by $\mathrm{FPWP}$ with different variances of communities size.
@@
In our final experiment, we examine the execution time of $\mathrm{FPWP}$ on different random networks. In particular, we use 5 random network models: a Forest Fire Network (FFN) [38], Barabási-Abert (BA) [39], Small world (SW) [40], Erdős-Rényi (ER) [41], and a clear community-structure network (CCS network). We use each of these models to generate three random networks, all with around 200 nodes and 600 edges, but different numbers of communities. To provide a community assignment for the computation of modularity, we use a modularity optimization on the original deterministic networks. We also control the number of communities, where $|C|=4$, $|C|=5$ and $|C|=6$ respectively.
@@
Figure 12 shows the running time of $\mathrm{FPWP}$ on networks with different network structures and numbers of communities. Generally, with an increasing number of communities the running time also increases, which fits the results in Figure 9. The visualization in Figure 11 suggests that a modularity optimization method may find a larger central community in FFN, which explains why the running time of FFN is always longer than for other networks. In fact, as shown in Figure 10, the larger the community size variance, the longer the running time.
@@
Real data
In this section, we apply $\mathrm{FPWP}$ and the sampling method on real networks. Notice that we do not use thresholding and weighting, because the experiments on synthetic data have clearly shown that these approaches should not be used to compute expected modularity. Figure 14 shows different network structures with different numbers of clusters. Here nodes of the same color are in the same community computed using a modularity maximization method, and the number of colors represents the number of communities. Figure 13 shows the execution time (vertical blue line) and the computed expected modularity (blue horizontal line) for our algorithm. Each red circle is the result of a different execution of the sampling method. For different executions, we have let the sampling method run for different amounts of time, so that we could inspect the balance between execution time and the likelihood of returning an accurate result for the sampling method. Figure 13 shows the same trends seen in the experiments with synthetic data. In these datasets, where the probability distributions are fixed and where clear community structures exist (we remind the reader that real data was obtained from high-modularity networks), the sampling method can produce accurate results in a short time, in accordance with our results on synthetic data.
@@
(a) Collaboration network $|E|=1224,|C|=2$
(b) PPI network $|E|=1185,|C|=5$
(c) Enron network $|E|=833,|C|=3$
(d) OSN
$|E|=1217,|C|=4$
@@
Discussion
In this paper, we evaluate alternative approaches for the calculation of expected modularity in probabilistic networks, including an original approach compatible with possible worlds semantics.
@@ Discussion
When compared with a brute-force approach, our new approach is several orders of magnitude faster, with $\mathrm{FPWP}$ reducing time complexity from exponential to polynomial without affecting accuracy. For a network with 25 edges, which is around the limit of what we could handle using exact computation methods before this paper, the execution time goes down from more than 15 hours (brute-force) to less than 5 seconds ($\mathrm{PWP}$) to a few microseconds ($\mathrm{FPWP}$). The execution time is however still high for large networks. In cases when the computation has to be repeated multiple times sequentially, for example inside unoptimized expected modularity maximization algorithms, one may have to consider using faster but less accurate approaches.
@@
We show that considering probabilities as weights, which can be a useful approach for other tasks, should not be used to compute expected modularity because it returns wrong results in general. The main reason is that node strengths (which equal node expected degrees in probabilistic networks) are multiplied in weighted modularity computation, while the expected product of degrees is not the same as the product of expected degrees in case of dependencies. Unfortunately, degrees in probabilistic networks are not independent, even when edge probabilities are, because edges are incident to two nodes and thus they either contribute to the degree of both or of none of them. A second reason is that weights are interpreted relative to each other in weighted modularity. So, for example, having all weights set to 0.01 or having them set to 1 makes no difference, while interpreting these numbers as probabilities means that in one case the network is probably empty, while in the second case all edges are almost certainly there. Identifying cases where using weighted modularity instead of expected modularity can still be appropriate is an interesting research question raised by this work.
@@
Thresholding is a common and very fast approach to handle probabilistic networks. In fact, we argue that even if the large majority of network studies currently use deterministic networks, these are in fact often uncertain networks where some more or less explicit decisions have been made about including or not edges based on the available information about the modeled system — that is, some thresholding has been performed. In this paper we show that thresholding is not a good approach for the task of computing expected modularity in general. However this is the case only when the network is almost deterministic. If we assume that all probabilities are at distance $\epsilon$ from 0 or 1, using thresholding we change edge probabilities from $1-\epsilon$ to 1 and from $\epsilon$ to 0, for all thresholds between $\epsilon$ and $1-\epsilon$. With small values of $\epsilon$, the change is small and the range of good thresholds is wide, making the method fast, accurate, and robust. However, as we are interested in handling probabilistic networks, it is not very useful to have a method that only works when the network is practically not probabilistic.
@@
While the amount of uncertainty has practically no effect on the accuracy and execution time of the new method we introduce in this paper, this is an important factor when sampling is used. This feature of sampling is already well acknowledged in the literature, where some methods have been proposed to tranform the original network into another probabilistic network of lower entropy [29, 42]. Sampling from a probabilistic network with probabilities very close to 0 or 1 immediately gives an accurate result, the same computed on the thresholded deterministic network. We have also noticed how the likelihood of producing an accurate value of expected modularity using sampling increases when the community structure is very clear, close to a set of cliques. This is due to the fact that modularity depends on the number of edges inside communities, and not on the position of these edges. As a result, when many (and most) edges are inside communities, many sampled networks will have a similar number of edges inside those communities, and thus similar values of modularity. Sampling from a narrow distribution has then a faster convergence. However, sampling can produce an inaccurate result, in addition to not giving any indication of whether the result is accurate or not. The more the network is uncertain, the less well sampling works.
@@
This paper opens the problem of how to use methods for the computation of expected modularity as a sub-procedure of a community detection method. Such an algorithm has not been proposed yet, despite the popularity of its deterministic counterpart. So far, clustering algorithms for probabilistic networks have been mostly developed outside of the network science research community, overlooking the popularity of the modularity objective function[16, 43, 44, 12]. This problem adds two interesting aspects to our work: first, when used as a heuristic inside an optimization algorithm, a less accurate calculation of modularity may still lead to the same clustering, so there is a question about when it is needed to be accurate and when a faster approximation is sufficient. That is, both sampling and considering probabilities as weights could be usable inside a community detection algorithm, even if their results may be inaccurate. Second, we may have to execute the modularity calculation many times, which poses an additional computational challenge in particular for large networks. These two aspects may allow the usage of different computational approaches[45, 46].
@@ Acknowledgments
This work has been partly funded by eSSENCE, an e-Science collaboration funded as a strategic research area of Sweden, by the Centre of Natural Hazards and Disaster Science (CNDS), by the Wallenberg AI, Autonomous Systems and Software Program WASP, and by the AI4Research initiative at Uppsala University.
Author contributions statement
Data Curation, Software, Investigation, Writing – Original Draft: X.S.; Methodology: X.S., M.M., C.R. Conceptualization, Writing – Review & Editing: X.S., M.M., C.R., F.S.; Supervision, Funding acquisition: M.M., C.R.
Data availability statement
The datasets generated and analysed during the current study are available in the repository: https://github.com/XINS3/Expected-modularity-calculation-over-uncertain-graph.
Additional information
The code to replicate the experiments is available here: https://github.com/XINS3/Expected-modularity-calculation-over-uncertain-graph; the authors declare no competing interests.